what is the amount of heat released when 94g of water at 80 degrees Celsius to form ice at -30 degrees Celsius
Answer #1

no of moles of H2O   = W/G.M.Wt

                               = 94/18   = 5.22moles

q1 = 94*4.18*(80-0)   = 31433.6J    = 31.4336KJ

q2 = n*H fus

     = 5.22mole*6.01KJ/mole = 31.37KJ

q3 = mcT

     = 94*2.09*30   = 5893.8J   = 5.8938KJ

total heat release q = q1 + q2 + q3

= 31.4336+31.37+5.8938   = 68.7KJ